# Zero-sum partitions of Abelian groups and their applications to magic-type labelings

### v3

@@ Abstract.
$|I(\Gamma)|\neq 1$, where $I(\Gamma)$ is the set of involutions in $\Gamma$, is a necessary condition for the existence of zero-sum partitions. In this paper, we show that the additional condition of $m_{i}\geq 4$ for every $1\leq i\leq t$, is sufficient. Moreover, we present some applications of zero-sum partitions to magic-type labelings of graphs.
@@
Abelian group, zero-sum sets, irregular labeling, magic-type labeling, anti-magic-type labeling
@@
This work was supported by program „Excellence initiative – research university” for the AGH University.
@@ 1.1. Preliminaries
For standard terms and notation in graph theory, the reader is referred to the textbook by Diestel [10] and the monograph by Brandstädt et al. [3]. An introduction to magic- and anti-magic-type labelings can be found in the monograph by Bača et al. [2], and the fundamental results in abstract algebra that we use can be found in the textbook by Gallian [13].
@@ 1.2. Main Problem
Let $\Gamma$ be an Abelian group such that $\Gamma\cong L\times H$, where $|L|=2^{\eta}$ and $|H|=\rho$ for a positive integer $\eta$ and an odd positive integer $\rho$, with $\rho>1$. The main contribution of this paper is to show that $\Gamma$ has $3$-ZSPP if $(\rho\bmod{6})\in\{1,3\}$. Moreover, for $(\rho\bmod{6})=5$, we show that $\Gamma$ has $4$-ZSPP. Therefore we improve the result from [8] and, towards proving Conjecture 1.9, we leave open only the question if, for $(\rho\bmod{6})=5$, $\Gamma$ has not only $4$-ZSPP, but also $3$-ZSPP. We also show that Conjecture 1.7 holds only for a group $\Gamma$ with $|I(\Gamma)|=3$ or $\Gamma\cong(\mathbb{Z}_{2})^{n}$ with any $n>1$. As a complement, we present some applications in irregular, magic-type, and anti-magic-type graph labeling to illustrate the relations of zero-sum group partitioning to graph theory.
@@ 2. Main Result
Assume now that $\beta=0$. If $|L^{*}|>7$, then, for $\alpha\geq 2$, the triple $(\alpha-2,0,l-\alpha+1)$ is realizable in $M^{*}$ and $(2,0,\gamma+\alpha-l-1)$ is realizable in $W$. For $\alpha\leq 1$, there is $r_{l-1}=5$, and $(\alpha+3,0,l-\alpha-2)$ is realizable in $M^{*}$ and $(1,2,\gamma+\alpha-l-2)$ is realizable in $W$. Observe that, in this case, by the construction of $W$, the two zero-sum $4$-subsets in $W$ can be split into four zero-sum $2$-subsets and we obtain a realization of $(\alpha,\beta,\gamma)$ in $\Gamma^{*}$. If $|L^{*}|=7$, then $\alpha=1$ and $L\cong(\mathbb{Z}_{2})^{3}$. Which implies that $|\Gamma|=8|H|=4+5\gamma$. Therefore, $\Gamma\cong(\mathbb{Z}_{2})^{3}\times H$ with $|H|\geq 13$, $\gamma=(|\Gamma|-4)/5=(8|H|-4)/5\geq 20$. Take the following three zero-sum $5$-subsets:
@@
By Theorem 1.6, we can assume that $|I(L)|\geq 7$. By Theorem 1.10, $L$ has $3$-ZSPP and there are $\phi,\varphi\in$Bij$(L)$ such that $a+\phi(a)+\varphi(a)=0$ for every $a\in L$ and $\phi(0)=\varphi(0)=0$ by fLemma 2.1.
@@
Assume first that $\Gamma\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$. By Theorem 1.6, $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{3}$ has $2$-ZSPP. The triples $(\alpha,\beta,\gamma)$ with $\beta<3$ were analyzed by a computer program we created, sample realizations can be found in the annexes. Therefore, for $A=(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{3}$ and $n_{1}=1$, by Lemma 2.3, we obtain that every triple $(\alpha,\beta,\gamma)$ is realizable in $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$.
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}(\{e_{1}+g,e_{2}+\phi(g),e_{1}+e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}\right).\end{array}$
The sequence $(|B|-1,0,0)$ is realizable in $\cup_{g\in B^{*}}\{a+g,\phi(g),-a+\varphi(g)\}$ and $(|B|,0,0)$ is realizable in $\cup_{g\in B}\{e_{1}+g,e_{2}+\phi(g),e_{1}+e_{2}+\varphi(g)\}$. Since $|B|-1+|B|=|L|/2-1\leq\alpha+\gamma$, we can find the remaining $\alpha+\gamma-|L|/2+1$ zero-sum $3$-subsets in
@@
Assume first that $\Gamma\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$. By Theorem 1.6, $\Gamma\cong(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{5}$ has $2$-ZSPP. The triples $(\alpha,\beta,\gamma)$ with $\beta<5$ were analyzed by a computer program we created, sample realizations can be found in the annexes. Therefore, for $A=(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{5}$ and $n_{1}=1$, by Lemma 2.3, we obtain that every triple $(\alpha,\beta,\gamma)$ is realizable in $\Gamma\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$.
@@
Every finite Abelian group $\Gamma$ has the $4$-Zero-Sum Partition Property if and only if $|I(\Gamma)|\neq 1$.
@@
Assume first that $e$ is even. Let $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},\ldots,m_{|R|/2+|I(\Gamma)|/3-2}=3$, $m_{|R|/2+|I(\Gamma)|/3-1},$ $m_{|R|/2+|I(\Gamma)|/3}=4$. Note that $\sum_{i=1}^{|R|/2+|I(\Gamma)|/3}m_{i}=|\Gamma^{*}|$. Suppose that there is a subset partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$ such that $|S_{i}|=m_{i}$ and $\sum_{s\in S_{i}}s=0$ for all $i$. Then $R\setminus\left(\bigcup_{i=1}^{|R|/2-1}S_{i}\right)=\{a,-a\}$ for some $a\in R$. Thus $a\in S_{j}$ for some $j\in\{|R|/2,\ldots,|R|/2+|I(\Gamma)|/3\}$ but then $\sum_{s\in S_{j}}s\neq 0$ since $I(\Gamma)\cup\{0\}$ is a subgroup of $\Gamma$. Thus for any $\iota_{1},\iota_{2},\iota_{3}\in I(\Gamma)$ there is $\iota_{1}+\iota_{2}\in I(\Gamma)$, $\iota_{1}+\iota_{2}+\iota_{3}\in I(\Gamma)\cup\{0\}$ but $a\notin I(\Gamma)\cup\{0\}$, a contradiction.
For $e$ odd we can show analogously that there is no partition of $\Gamma^{*}$ for $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},$ $\ldots$,… , $m_{|R|/2+(|I(\Gamma)|+2)/3}=3$.∎
@@ 3. Some Applications
In this section, we present some applications of Corollary 2.6 to magic-type labelings. Generally speaking, such a labeling of a graph $G=(V,E)$ is a mapping from only $V$ or $E$, or their union $V\cup E$, to a set of labels, which most often is a set of integers or elements of a group. Then the weight of a graph element is typically the sum of labels of the adjacent or incident elements of one or both types. When the weight of all elements is required to be equal, then we speak of a magic-type labeling; when the weights should be all different, then we speak of an anti-magic-type labeling.
3.1. Group irregular labeling for digraphs
@@
3.2. Magic and anti-magic labeling in graphs with twins
Froncek[12] defined the notion of a group distance magic labeling of a graph $G=(V,E)$. In such a labeling, the vertices of the graph are labeled through a bijection with the elements of an Abelian group $\Gamma$. The weight of each vertex is computed as the sum (in $\Gamma$) of the labels assigned to its neighbors. If all weights are the same, then it is a $\Gamma$-distance magic labeling. If all weights are different, then it is a $\Gamma$-distance anti-magic labeling [1].
For ease of presentation of the results in this subsection we will create labelings using only the non-zero elements of $\Gamma$, and speak of $\Gamma^{*}$-distance magic labeling, etc. It is easy to see that these results can be amended to results on labelings with all elements of $\Gamma$ by dealing with the 00 element accordingly.
@@ 3.2. Magic and anti-magic labeling in graphs with twins
Based on the above observations, we can create $\Gamma^{*}$-distance magic and anti-magic labelings for groups and graphs that satisfy certain conditions detailed in the following propositions.
@@
Let $G=(V,E)$ be a graph with a partition of the vertex set into subsets of false twins $\{V_{i}\}_{i=1}^{t}$ for a positive integer $t$. Let $m_{i}=|V_{i}|$ for every $1\leq i\leq t$. Let $\Gamma$ be an Abelian group of order $m$ such that $\sum_{i=1}^{t}m_{i}=m-1$. If the non-zero elements of $\Gamma$ can be partitioned into zero-sum subsets $\{S_{i}\}_{i=1}^{t}$ such that $|S_{i}|=m_{i}$ for every $1\leq i\leq t$, then $G$ has a $\Gamma^{*}$-distance magic labeling.
@@
Let $G=(V,E)$ be a graph with a partition of the vertex set into subsets of true twins $\{V_{i}\}_{i=1}^{t}$ for a positive integer $t$. Let $m_{i}=|V_{i}|$ for every $1\leq i\leq t$. Let $\Gamma$ be an Abelian group of order $m$ such that $\sum_{i=1}^{t}m_{i}=m-1$. If the non-zero elements of $\Gamma$ can be partitioned into zero-sum subsets $\{S_{i}\}_{i=1}^{t}$ such that $|S_{i}|=m_{i}$ for every $1\leq i\leq t$, then $G$ has a $\Gamma^{*}$-distance anti-magic labeling.
@@
The following corollaries are direct consequences of Propositions 3.7 and 3.8, Corollary 2.6, and the properties of twins mentioned above. Note that we say that a vertex has $k$ twins if it has $k$ twins different from itself.
@@
Let $G=(V,E)$ be a graph in which every vertex has at least $3$ false twins. Let $\Gamma$ be a finite Abelian group of order $|V|+1$ such that $|I(\Gamma)|\neq 1$. Then $G$ has a $\Gamma^{*}$-distance magic labeling.
Notice that Corollary 3.9 extends the results on $\Gamma$-distance magic labelings of complete $k$-partite graphs presented in [7].
@@
Let $G=(V,E)$ be a graph in which every vertex has at least $3$ true twins. Let $\Gamma$ be a finite Abelian group of order $|V|+1$ such that $|I(\Gamma)|\neq 1$. Then $G$ has a $\Gamma^{*}$-distance anti-magic labeling.
@@
Recall that $X$ is an empty graph if $E(X)=\emptyset$. In other words, its vertex set is stable. We use $K_{k}$ to denote the complete graph on $k$ vertices and $\overline{K_{k}}$ for the empty graph on $k$ vertices. The following corollaries are direct consequences of Corollaries 3.9 and 3.10.
@@
Let $G=(V,E)$ be any graph. Then there exists a $G$-join $H$ of a sequence of empty graphs of appropriate orders such that every vertex in $H$ has at least $3$ false twins. In particular, $H$ has a $\Gamma^{*}$-distance magic labeling for every finite Abelian group $\Gamma$ with $|\Gamma|=|V(H)|+1$ and $|I(\Gamma)|\neq 1$.
@@
Let $G=(V,E)$ be any graph. Then there exists a $G$-join $H$ of a sequence of complete graphs of appropriate orders such that every vertex in $H$ has at least $3$ true twins. In particular, $H$ has a $\Gamma^{*}$-distance anti-magic labeling for every finite Abelian group $\Gamma$ with $|\Gamma|=|V(H)|+1$ and $|I(\Gamma)|\neq 1$.

### v4

# Zero-sum partitions of Abelian groups and their applications to magic- and antimagic-type labelings

@@ Abstract.
$|I(\Gamma)|\neq 1$, where $I(\Gamma)$ is the set of involutions in $\Gamma$, is a necessary condition for the existence of zero-sum partitions. In this paper, we show that the additional condition of $m_{i}\geq 4$ for every $1\leq i\leq t$, is sufficient. Moreover, we present some applications of zero-sum partitions to magic- and antimagic-type labelings of graphs.
@@
Abelian group, zero-sum sets, irregular labeling, magic-type labeling, antimagic-type labeling
@@
This work was supported by program “Excellence initiative – research university” for the AGH University.
@@ 1.1. Preliminaries
For standard terms and notation in graph theory, the reader is referred to the textbook by Diestel [10] and the monograph by Brandstädt et al. [3], an introduction to magic- and antimagic-type labelings can be found in the monograph by Bača et al. [2], and the fundamental results in abstract algebra that we use can be found in the textbook by Gallian [13].
@@ 1.2. Main Problem
Let $\Gamma$ be an Abelian group such that $\Gamma\cong L\times H$, where $|L|=2^{\eta}$ and $|H|=\rho$ for a positive integer $\eta$ and an odd positive integer $\rho$, with $\rho>1$. The main contribution of this paper is to show that $\Gamma$ has $3$-ZSPP if $(\rho\bmod{6})\in\{1,3\}$. Moreover, for $(\rho\bmod{6})=5$, we show that $\Gamma$ has $4$-ZSPP. Therefore we improve the result from [8] and, towards proving Conjecture 1.9, we leave open only the question if, for $(\rho\bmod{6})=5$, $\Gamma$ has not only $4$-ZSPP, but also $3$-ZSPP. We also show that Conjecture 1.7 holds only for a group $\Gamma$ with $|I(\Gamma)|=3$ or $\Gamma\cong(\mathbb{Z}_{2})^{n}$ with any $n>1$. As a complement, we present some applications in irregular, magic-type, and antimagic-type graph labeling to illustrate the relations of zero-sum group partitioning to graph theory.
@@ 2. Main Result
Assume now that $\beta=0$. If $|L^{*}|>7$, then, for $\alpha\geq 2$, the triple $(\alpha-2,0,l-\alpha+1)$ is realizable in $M^{*}$ and $(2,0,\gamma+\alpha-l-1)$ is realizable in $W$. For $\alpha\leq 1$, there is $r_{l-1}=5$, and $(\alpha+3,0,l-\alpha-2)$ is realizable in $M^{*}$ and $(1,2,\gamma+\alpha-l-2)$ is realizable in $W$. Observe that, in this case, by the construction of $W$, the two zero-sum $4$-subsets in $W$ can be split into four zero-sum $2$-subsets and we obtain a realization of $(\alpha,\beta,\gamma)$ in $\Gamma^{*}$. If $|L^{*}|=7$, then $\alpha=1$ and $L\cong(\mathbb{Z}_{2})^{3}$, which implies that $|\Gamma|=8|H|=4+5\gamma$. Therefore, $\Gamma\cong(\mathbb{Z}_{2})^{3}\times H$ with $|H|\geq 13$, $\gamma=(|\Gamma|-4)/5=(8|H|-4)/5\geq 20$. Take the following three zero-sum $5$-subsets:
@@
By Theorem 1.6, we can assume that $|I(L)|\geq 7$. By Theorem 1.10, $L$ has $3$-ZSPP and there are $\phi,\varphi\in$Bij$(L)$ such that $a+\phi(a)+\varphi(a)=0$ for every $a\in L$ and $\phi(0)=\varphi(0)=0$ by Lemma 2.1.
@@
Assume first that $\Gamma^{*}\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$. By Theorem 1.6, $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{3}$ has $2$-ZSPP. The triples $(\alpha,\beta,\gamma)$ with $\beta<3$ were analyzed by a computer program we created, sample realizations can be found in the annexes. Therefore, for $A=(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{3}$ and $n_{1}=1$, by Lemma 2.3, we obtain that every triple $(\alpha,\beta,\gamma)$ is realizable in $\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{3}\right)^{*}$.
@@
$\begin{array}[]{ll}\Gamma^{*}=&\{a,-a\}\cup\left(\bigcup\limits_{\begin{subarray}{c}g\in B^{*}\end{subarray}}\{a+g,\phi(g),-a+\varphi(g)\}\cup\bigcup\limits_{\begin{subarray}{c}g\in B\end{subarray}}(\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}\right)\\ &\cup\left(\bigcup\limits_{g\in B}\left\{c+g,d+\phi(g),-c-d+\varphi(g),-c-g,-d-\phi(g),c+d-\varphi(g)\right\}\right).\end{array}$
The sequence $(|B|-1,0,0)$ is realizable in $\cup_{g\in B^{*}}\{a+g,\phi(g),-a+\varphi(g)\}$ and $(|B|,0,0)$ is realizable in $\cup_{g\in B}\{e_{1}+g,e_{2}+\phi(g),-e_{1}-e_{2}+\varphi(g)\}$. Since $|B|-1+|B|=|L|/2-1\leq\alpha+\gamma$, we can find the remaining $\alpha+\gamma-|L|/2+1$ zero-sum $3$-subsets in
@@
Assume first that $\Gamma\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$. By Theorem 1.6, $\Gamma\cong(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{5}$ has $2$-ZSPP. The triples $(\alpha,\beta,\gamma)$ with $\beta<5$ were analyzed by a computer program we created, sample realizations can be found in the annexes. Therefore, for $A=(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{5}$ and $n_{1}=1$, by Lemma 2.3, we obtain that every triple $(\alpha,\beta,\gamma)$ is realizable in $\Gamma^{*}\cong\left((\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{5}\right)^{*}$.
@@
A finite Abelian group $\Gamma$ has the $4$-Zero-Sum Partition Property if and only if $|I(\Gamma)|\neq 1$.
@@
Assume first that $e$ is even. Let $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},\ldots,m_{|R|/2+|I(\Gamma)|/3-3}=3$, $m_{|R|/2+|I(\Gamma)|/3-2},$ $m_{|R|/2+|I(\Gamma)|/3-1}=4$. Note that $\sum_{i=1}^{|R|/2+|I(\Gamma)|/3-1}m_{i}=|\Gamma^{*}|$. Suppose that there is a subset partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$ such that $|S_{i}|=m_{i}$ and $\sum_{s\in S_{i}}s=0$ for all $i$. Then $R\setminus\left(\bigcup_{i=1}^{|R|/2-1}S_{i}\right)=\{a,-a\}$ for some $a\in R$. Thus $a\in S_{j}$ for some $j\in\{|R|/2,\ldots,|R|/2+|I(\Gamma)|/3-1\}$ but then $\sum_{s\in S_{j}}s\neq 0$ since $I(\Gamma)\cup\{0\}$ is a subgroup of $\Gamma$. Thus for any $\iota_{1},\iota_{2},\iota_{3}\in I(\Gamma)$ there is $\iota_{1}+\iota_{2}\in I(\Gamma)$, $\iota_{1}+\iota_{2}+\iota_{3}\in I(\Gamma)\cup\{0\}$ but $a\notin I(\Gamma)\cup\{0\}$, a contradiction.
For $e$ odd we can show analogously that there is no partition of $\Gamma^{*}$ for $m_{1},\dots,m_{|R|/2-1}=2$, $m_{|R|/2},$ $\ldots$,… , $m_{|R|/2+(|I(\Gamma)|+2)/3-1}=3$.∎
@@ 3. Some Applications
In this section, we present some applications of Corollary 2.6 to magic- and antimagic-type labelings. Generally speaking, such a labeling of a graph $G=(V,E)$ is a mapping from only $V$ or $E$, or their union $V\cup E$, to a set of labels, which most often is a set of integers or elements of a group. Then the weight of a graph element is typically the sum of labels of the adjacent or incident elements of one or both types. When the weight of all elements is required to be equal, then we speak of a magic-type labeling; when the weights should be all different, then we speak of an antimagic-type labeling.
3.1. $\Gamma$-irregular labeling for digraphs
@@
3.2. $\Gamma$-antimagic labeling
Let $\Gamma$ be an Abelian group. A $\Gamma$-antimagic labeling of a graph $G$ is defined as a bijection $f\colon E(G)\to\Gamma$, where the weight of each vertex, which is the sum of the labels on its incident edges, is distinct for all vertices. Kaplan et al. [16] demonstrated that if $\Gamma$ has a unique involution, then no tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling. They conjectured that a tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling if and only if $\Gamma$ does not have a unique involution. By applying the same approach used in [16] for 2-trees, and using Corollary 2.6, we derive the following result:
Corollary 3.7.
Every $4$-tree [^1] A $4$-tree $T$ is a rooted tree, where every vertex that is not a leaf has at least four children. $T$ of order $|\Gamma|$ admits a $\Gamma$-antimagic labeling if and only if $|I(\Gamma)|\neq 1$.
Proof.
The necessity of the condition follows directly from the work of Kaplan et al. (see Theorem E, [16]). Thus, let us assume that $|I(\Gamma)|\neq 1$. Let $\{v_{i}\}_{i=1}^{t}$ be the vertices of $T$ which are not leaves. Let us denote their corresponding numbers of children by $\{m_{i}\}_{i=1}^{t}$. Thus $\sum_{i=1}^{t}m_{i}=|\Gamma|-1$. Since $T$ is a $4$-tree, we have that $m_{i}\geq 4$ for every $1\leq i\leq t$. The group $\Gamma$ has $4$-ZSPP by Corollary 2.6, and there exists a zero-sum partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$. For every $1\leq i\leq t$, we label the edges of the set $E_{i}=\{v_{i}w\colon w$ is a child of $v_{i}\}$ by the elements of $S_{i}$. The edges of $T$ are labeled bijectively with the non-zero elements of $\Gamma$, and the sum of the labels in every $E_{i}$ is 00. Since every vertex of $T$, except the root (whose weight is 00), has a unique parent, it easily follows that the weights are pairwise distinct. ∎
3.3. $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Froncek [12] defined the notion of a group-distance-magic labeling of a graph $G=(V,E)$. In such a labeling, the vertices of the graph are labeled through a bijection with the elements of an Abelian group $\Gamma$. The weight of each vertex is computed as the sum (in $\Gamma$) of the labels assigned to its neighbors. If all weights are the same, then it is a $\Gamma$-distance-magic labeling. If all weights are different, then it is a $\Gamma$-distance-antimagic labeling [1].
For ease of presentation of the results in this subsection we will create labelings using only the non-zero elements of $\Gamma$, and speak of $\Gamma^{*}$-distance-magic labeling, etc. It is easy to see that these results can be amended to results on labelings with all elements of $\Gamma$ by dealing with the 00 element accordingly.
@@ 3.3. $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Based on the above observations, we can create $\Gamma^{*}$-distance-magic and -antimagic labelings for groups and graphs that satisfy certain conditions detailed in the following propositions.
@@
Let $G=(V,E)$ be a graph with a partition of the vertex set into subsets of false twins $\{V_{i}\}_{i=1}^{t}$ for a positive integer $t$. Let $m_{i}=|V_{i}|$ for every $1\leq i\leq t$. Let $\Gamma$ be an Abelian group of order $m$ such that $\sum_{i=1}^{t}m_{i}=m-1$. If the non-zero elements of $\Gamma$ can be partitioned into zero-sum subsets $\{S_{i}\}_{i=1}^{t}$ such that $|S_{i}|=m_{i}$ for every $1\leq i\leq t$, then $G$ has a $\Gamma^{*}$-distance-magic labeling.
@@
Let $G=(V,E)$ be a graph with a partition of the vertex set into subsets of true twins $\{V_{i}\}_{i=1}^{t}$ for a positive integer $t$. Let $m_{i}=|V_{i}|$ for every $1\leq i\leq t$. Let $\Gamma$ be an Abelian group of order $m$ such that $\sum_{i=1}^{t}m_{i}=m-1$. If the non-zero elements of $\Gamma$ can be partitioned into zero-sum subsets $\{S_{i}\}_{i=1}^{t}$ such that $|S_{i}|=m_{i}$ for every $1\leq i\leq t$, then $G$ has a $\Gamma^{*}$-distance-antimagic labeling.
@@
The following corollaries are direct consequences of Propositions 3.8 and 3.9, Corollary 2.6, and the properties of twins mentioned above. Note that we say that a vertex has $k$ twins if it has $k$ twins different from itself.
@@
Let $G=(V,E)$ be a graph in which every vertex has at least $3$ false twins. Let $\Gamma$ be a finite Abelian group of order $|V|+1$ such that $|I(\Gamma)|\neq 1$. Then $G$ has a $\Gamma^{*}$-distance-magic labeling.
Notice that Corollary 3.10 extends the results on $\Gamma$-distance-magic labelings of complete $k$-partite graphs presented in [7].
@@
Let $G=(V,E)$ be a graph in which every vertex has at least $3$ true twins. Let $\Gamma$ be a finite Abelian group of order $|V|+1$ such that $|I(\Gamma)|\neq 1$. Then $G$ has a $\Gamma^{*}$-distance-antimagic labeling.
@@
Recall that $X$ is an empty graph if $E(X)=\emptyset$. In other words, its vertex set is stable. We use $K_{k}$ to denote the complete graph on $k$ vertices and $\overline{K_{k}}$ for the empty graph on $k$ vertices. The following corollaries are direct consequences of Corollaries 3.10 and 3.11.
@@
Let $G=(V,E)$ be any graph. Then there exists a $G$-join $H$ of a sequence of empty graphs of appropriate orders such that every vertex in $H$ has at least $3$ false twins. In particular, $H$ has a $\Gamma^{*}$-distance-magic labeling for every finite Abelian group $\Gamma$ with $|\Gamma|=|V(H)|+1$ and $|I(\Gamma)|\neq 1$.
@@
Let $G=(V,E)$ be any graph. Then there exists a $G$-join $H$ of a sequence of complete graphs of appropriate orders such that every vertex in $H$ has at least $3$ true twins. In particular, $H$ has a $\Gamma^{*}$-distance-antimagic labeling for every finite Abelian group $\Gamma$ with $|\Gamma|=|V(H)|+1$ and $|I(\Gamma)|\neq 1$.